# Understanding High-Dimensional Bayesian Optimization

### v1

@@ Abstract
Recent work reported that simple Bayesian optimization methods perform well for high-dimensional real-world tasks, seemingly contradicting prior work and tribal knowledge. This paper investigates the ‘why’. We identify fundamental challenges that arise in high-dimensional Bayesian optimization and explain why recent methods succeed. Our analysis shows that vanishing gradients caused by Gaussian process initialization schemes play a major role in the failures of high-dimensional Bayesian optimization and that methods that promote local search behaviors are better suited for the task. We find that maximum likelihood estimation of Gaussian process length scales suffices for state-of-the-art performance. Based on this, we propose a simple variant of maximum likelihood estimation called MSR that leverages these findings to achieve state-of-the-art performance on a comprehensive set of real-world applications. We also present targeted experiments to illustrate and confirm our findings.
@@ 1 Introduction
Bayesian optimization (BO) has found wide-spread adoption for optimizing expensive-to-evaluate black-box functions that appear in aerospace engineering (Lukaczyk et al., 2014; Lam et al., 2018), drug discovery (Negoescu et al., 2011), robotics (Lizotte et al., 2007; Calandra et al., 2016; Rai et al., 2018; Mayr et al., 2022) or finance (Baudiš & Pošík, 2014).
@@
While BO has proven reliable in low-dimensional settings, high-dimensional spaces are challenging due to the curse of dimensionality (COD) that demands exponentially more data points to maintain the same precision with increasing problem dimensionality. Several approaches have extended BO to high-dimensional spaces under additional assumptions on the objective function that lower the data demand, such as additivity (Duvenaud et al., 2011; Kandasamy et al., 2015; Han et al., 2021; Ziomek & Ammar, 2023) or the existence of a low-dimensional active subspace (Wang et al., 2016; Nayebi et al., 2019; Letham et al., 2020; Papenmeier et al., 2022). Without such assumptions, it was widely believed that BO with a Gaussian process (GP) surrogate is limited to approximately 20 dimensions for common evaluation budgets (Frazier, 2018; Moriconi et al., 2020). Recently, Hvarfner et al. (2024) and Xu & Zhe (2024) reported that simple BO methods perform well on high-dimensional real-world benchmarks, often surpassing the performance of more sophisticated algorithms.
@@
Due to its many impactful applications, high-dimensional Bayesian optimization (HDBO) has seen active research in recent years (Binois & Wycoff, 2022; Papenmeier et al., 2023). While the boundaries have been pushed significantly, the causes of performance gains have not always been thoroughly identified. For example, in the case of the BODi (Deshwal et al., 2023) and COMBO (Oh et al., 2019) algorithms, later work found that the methods benefited from specific benchmark structures prevalent in their evaluation (Papenmeier et al., 2023). Similarly, an evaluation of Nayebi et al. (2019) showed that the performance of some prior methods is sensitive to the location of the optimum in the search space.
@@
Thus, answering the recent call for more scrutiny and exploratory research (Herrmann et al., 2024), we first identify fundamental challenges arising in high dimensions and then examine state-of-the-art HDBO methods to understand how they mitigate these obstacles. Equipped with these insights, we propose a simpler approach that uses maximum likelihood estimation (MLE) of the GP length scales called MLE Scaled with RAASP (MSR). We demonstrate that MSR is sufficient for state-of-the-art HDBO performance without the need for specifying a prior belief on length scales as in maximum a-posteriori estimation (MAP). We note that practitioners usually do not possess such priors and instead rely on empirical performances on benchmarks.In particular, we change the initialization of length scales to avoid vanishing gradients of the GP likelihood function that easily occur in high-dimensional spaces but, so far, have been overlooked for BO. Furthermore, we provide empirical evidence suggesting that good BO performance on extremely high-dimensional problems (on the order of $1000$ dimensions) is due to local search behavior and not to a well-fit surrogate model. In summary, we make the following contributions.
@@
    We identify fundamental challenges that arise in high-dimensional Bayesian optimization and explain why recent methods succeed. We show that vanishing gradients and local search behaviors are important in HDBO.
@@ 3 Facets of the Curse of Dimensionality
This section discusses how the curse of dimensionality impacts high-dimensional Bayesian optimization and techniques to mitigate these challenges.
@@ 3.1 Vanishing Gradients at Model Fitting
While these methods mitigate the issue of vanishing gradients, increasing length scales by any constant factor does not always solve it, as Fig. 1 indicates. Whether scaling the hyperprior of the length scale with the problem’s dimensionality achieves a good fit of the surrogate model depends on the properties of $f$. The success of the two methods described above is related to the effect of the increased length scales on mitigating the problem of vanishing gradients. If the underlying function varies quickly, the GP needs to use a short length scale to model $f$. In such cases, the GP cannot model the function globally, as shown in Appendix B.4.
@@ 3.2 Vanishing Gradients of the Acquisition Function
Differing slightly from TuRBO’s approach, RAASP sampling has been implemented in BoTorch’s AF maximization and can optionally be enabled with the parameter sample_around_best. BoTorch augments a set of globally sampled candidates with the RAASP samples, resulting in twice as many initial candidates. It perturbs the best-performing points by replacing the value of each dimension with a probability of $\min\left(1,\frac{20}{d}\right)$ by a value drawn from a Gaussian distribution, truncated to stay within the bounds of the search space. BoTorch then chooses the points of maximum initial acquisition value to start the GD optimization of the acquisition function.
With increasing dimensionality or descending length scale, the starting points for the multi-start GD routine chosen by the AF maximizer are increasingly more likely to originate from the RAASP samples. Fig. 4 (right panel) illustrates this. Here, we draw realizations from GPs, initialized with different dimensionalities ($x$-axis) and length scales ($y$-axis). For each realization, we maximize the acquisition function with RAASP sampling and plot the percentage of candidates of maximum acquisition value originating from the RAASP samples across all candidates. A higher percentage indicates a more ‘local’ sampling. We further average across five random restarts. The percentage of RAASP candidates with maximum acquisition value increases with the input dimensionality and decreases with the length scale.
@@
In general, when the length scales of the GP are short and the dimensionality is high, BO shows a local-search-like behavior with RAASP sampling and a random search behavior without it. We demonstrate this using the observation traveling salesperson distance (OTSD(Papenmeier et al., 2025), which quantifies an algorithm’s exploration by finding the shortest path connecting all observations made up to a certain iteration. The OTSD curve of an algorithm $A$ consistently lying above the curve of algorithm $B$ indicates that algorithm $A$ is more explorative as its observations are more spread more evenly across $\mathcal{X}$. The OTSD is always monotonic increasing. Fig. 5 shows the OTSDs for $100$-dimensional GPs with different length scales, each initialized with $10$ random samples in the design of experiments (DOE) phase and subsequently optimized with LogEI and RAASP sampling for $20$ iterations. Unless the model length scale is sufficiently long for the AF gradient not to vanish (as for $\ell=0.5$ in the right panel of Fig. 5), the AF maximizer picks one of the initial random candidates without further optimizing it. This is supported by the trajectories for the BO phase (iteration $\geq 9$) following the trajectory of the DOE phase (iteration $<9$) for $\ell=0.05$ and $\ell=0.28$ in the right panel of Fig. 5. We generally recommend the RAASP sampling method as it improves BO by automatically reverting to local search when encountering flat AFs.
@@ 4 Discussion
We propose a simple initialization for the gradient-based optimizer used for fitting the length scales of the GP surrogate via MLE and evaluate its performance for BO tasks. In what follows, we suppose a BO algorithm with a $\nicefrac{{5}}{{2}}$-ARD-Matérn kernel and LogEI (Ament et al., 2024). To address the issue of vanishing gradients at the start of the MLE optimization, we choose the initial length scale as 0.1 and scale with $\sqrt{d}$ to account for the increasing distances of the randomly sampled DOE points. Thus, the initial length scale used in the optimization is $\frac{\sqrt{d}}{10}$, and we refer to this new Bayesian optimization method as ‘MLE scaled’. The second method in the evaluation is the same BO method, but now using the default value $\ell=\ln 2$ of GPyTorch as initial length scale in MLE. This value is not scaled with $d$. Based on our analysis of the impact of RAASP sampling when optimizing the AF, we combine ‘MLE scaled’ with RAASP sampling and call the method ‘MLE Scaled with RAASP’ (MSR). The fifth method DSP is the new default in BoTorch (Balandat et al., 2020), which uses a MAP estimate of length scales and initializes the optimization with the mode of the dimensionality-scaled length scale prior (DSP(Hvarfner et al., 2024). We also compare against SAASBO (Eriksson & Jankowiak, 2021) and Bounce (Papenmeier et al., 2023). SAASBO has a large computational runtime. Hence, we run it only for 500 iterations and terminate runs that exceed 72 hours. That is why SAASBO has fewer evaluations for Ant and Humanoid.
Our benchmarks are the $124$-dimensional soft-constrained version of the Mopta08 benchmark (Jones, 2008) introduced by Eriksson & Jankowiak (2021), the 180-dimensional Lasso-DNA benchmark (Šehić et al., 2022), as well as two 888- and 6392-dimensional Mujoco benchmarks used by Hvarfner et al. (2024). The first two benchmarks are noise-free, while the others exhibit observational noise.
@@
Fig. 8 shows the performance of the BO methods on the four real-world applications. Each plot gives the average objective value of the best solution found so far in terms of the number of iterations. We show the ranking of the methods according to the final performance in Table 1 in Appendix B.1. The confidence bands indicate the standard error of the mean. MSR achieves competitive performance across all benchmarks, matching the SOTA DSP. Notably, MSR outperforms DSP on $124d$ Mopta08 and $888d$ Ant, and performs slightly worse than DSP other benchmarks. Bounce (Papenmeier et al., 2023) is consistently outperformed by MSR, MLE, and DSP but surpasses SAASBO (Eriksson & Jankowiak, 2021), especially on Ant. Although the constant length scale initialization ($\ell=\ln 2$) without RAASP achieves satisfactory results on lower-dimensional benchmarks such as $124d$ Mopta08 and $180d$ Lasso-DNA, it fails on higher-dimensional benchmarks like $888d$ Ant and $6392d$ Humanoid. We attribute this breakdown to vanishing gradients as shown in Fig. 12 in Appendix B.2.
@@
For the 124-dimensional Mopta08 and 180-dimensional Lasso-DNA applications, ‘MLE ($\ell=\ln 2$)’ learns length scales similar to the other methods. However, this constant initialization strategy fails to make progress from the initial value for the more high-dimensional problems in the bottom row of Fig. 9. We attribute this to vanishing gradients of the MLL as discussed in Sec. 3 and highlighted in Fig. 12 in Appendix B.2.
@@
Fig. 9 shows a surprising behavior for DSP. As one would anticipate, the estimated length scales are typically close to the mode of the hyperprior at the start of the optimization. However, they then converge to an even higher value on all benchmarks but the $6392$-dimensional Humanoid benchmark. Furthermore, the deviation from the prior mode is more pronounced for the lower-dimensional benchmarks, being in line with our analysis of the bias-variance trade-off in Sec. 3.3. MLE with scaled initial length scales (‘MLE (scaled)’) initially has higher length scale values than all other methods. The resulting estimates vary significantly for the high-dimensional Ant and Humanoid problems, supporting our analysis in Sec. 3.3 where we study the comparatively high variance of MLE compared to MAP.
@@
In Fig. 9, all methods converge to similarly long length scales on the Mopta08 and Lasso-DNA benchmarks. This is likely attributable to a specific property of these popular benchmarks, as we discuss in Appendix C. In short, these benchmarks have a simple structure that benefits models with long length scales, posing the risk of algorithms being ‘overfitted’ to these benchmarks.
@@ 5 Conclusions and Future Work
Our analysis reveals fundamental challenges in HDBO while offering practical insights for improving HDBO methods. We demonstrate that common approaches for fitting Gaussian processes cause vanishing gradients in high dimensions. We propose an initialization for MLE that achieves state-of-the-art performance on established HDBO benchmarks without requiring the assumptions of maximum a-posteriori estimation. Finally, we provide empirical evidence that combining MLE with RAASP sampling reduces the variance of the length scale estimates and yields values closer to the ones learned by DSP, providing a fresh argument for the inclusion of RAASP sampling in HDBO.
In future work, we will continue to carefully vet popular benchmarks and propose novel, challenging benchmarks that preserve the traits of real-world applications. Furthermore, this work emphasizes the importance of numerical stability for the performance of Bayesian optimization in high dimensions. Thus, we propose approaching the development of models and acquisition functions from this perspective.
Impact.

### v2

@@ Abstract
Recent work reported that simple Bayesian optimization (BO) methods perform well for high-dimensional real-world tasks, seemingly contradicting prior work and tribal knowledge. This paper investigates why. We identify underlying challenges that arise in high-dimensional BO and explain why recent methods succeed. Our empirical analysis shows that vanishing gradients caused by Gaussian process (GP) initialization schemes play a major role in the failures of high-dimensional Bayesian optimization (HDBO) and that methods that promote local search behaviors are better suited for the task. We find that maximum likelihood estimation (MLE) of GP length scales suffices for state-of-the-art performance. Based on this, we propose a simple variant of MLE called MSR that leverages these findings to achieve state-of-the-art performance on a comprehensive set of real-world applications. We present targeted experiments to illustrate and confirm our findings.
@@ 1 Introduction
BO has found wide-spread adoption for optimizing expensive-to-evaluate black-box functions that appear in aerospace engineering (Lukaczyk et al., 2014; Lam et al., 2018), drug discovery (Negoescu et al., 2011), robotics (Lizotte et al., 2007; Calandra et al., 2016; Rai et al., 2018; Mayr et al., 2022) or finance (Baudiš & Pošík, 2014).
@@
While BO has proven reliable in low-dimensional settings, high-dimensional spaces are challenging due to the curse of dimensionality (COD) that demands exponentially more data points to maintain the same precision with increasing problem dimensionality. Several approaches have extended BO to high-dimensional spaces under additional assumptions on the objective function that lower the data demand, such as additivity (Duvenaud et al., 2011; Kandasamy et al., 2015; Hoang et al., 2018; Han et al., 2021; Ziomek & Ammar, 2023; Bardou et al., 2024) or the existence of a low-dimensional active subspace (Wang et al., 2016; Nayebi et al., 2019; Letham et al., 2020; Papenmeier et al., 2022). Without such assumptions, it was widely believed that BO with a GP surrogate is limited to approximately 20 dimensions for common evaluation budgets (Frazier, 2018; Moriconi et al., 2020). Recently, Hvarfner et al. (2024) and Xu & Zhe (2024) reported that simple BO methods perform well on high-dimensional real-world benchmarks, often surpassing the performance of more sophisticated algorithms.
@@
Due to its many impactful applications, HDBO has seen active research in recent years (Binois & Wycoff, 2022; Papenmeier et al., 2023). While the boundaries have been pushed significantly, the causes of performance gains have not always been thoroughly identified. For example, in the case of the BODi (Deshwal et al., 2023) and COMBO (Oh et al., 2019) algorithms, later work found that the methods benefited from specific benchmark structures prevalent in their evaluation (Papenmeier et al., 2023). Similarly, an evaluation of Nayebi et al. (2019) showed that the performance of some prior methods is sensitive to the location of the optimum in the search space.
@@
This paper is motivated by the recent call for more scrutiny and exploratory research (Herrmann et al., 2024). By exhaustive experimentation, we first identify underlying challenges arising in high dimensions and then examine state-of-the-art HDBO methods to understand how they mitigate these obstacles. Equipped with these insights, we propose a simpler approach that uses MLE of the GP length scales called MLE Scaled with RAASP (MSR). We demonstrate that MSR is sufficient for state-of-the-art HDBO performance without the need for specifying a prior belief on length scales as in maximum a-posteriori estimation (MAP). We note that practitioners usually do not possess such priors and instead rely on empirical performances on benchmarks.In particular, we change the initialization of length scales to avoid vanishing gradients of the GP likelihood function that easily occur in high-dimensional spaces but, so far, have been overlooked for BO. Furthermore, we provide empirical evidence suggesting that good BO performance on extremely high-dimensional problems (on the order of $1000$ dimensions) is due to local search behavior and not to a well-fit surrogate model. In summary, we make the following contributions.
@@
    We identify underlying challenges that arise in HDBO and explain why recent methods succeed. We show that vanishing gradients and local search behaviors are important in HDBO.
@@ 3 Facets of the Curse of Dimensionality
This section discusses how the curse of dimensionality impacts HDBO and techniques to mitigate these challenges.
@@ 3.1 Vanishing Gradients at Model Fitting
While these methods mitigate the issue of vanishing gradients, increasing length scales by any constant factor does not always solve it, as Fig. 1 indicates. Whether scaling the hyperprior of the length scale with the problem’s dimensionality achieves a good fit of the surrogate model depends on the properties of $f$. The success of the two methods described above is related to the effect of the increased length scales on mitigating the problem of vanishing gradients. If the underlying function varies quickly, the GP needs to use a short length scale to model $f$. In such cases, the GP cannot model the function globally, as shown in Appendix C.4.
@@ 3.2 Vanishing Gradients of the Acquisition Function
Differing slightly from TuRBO’s approach, RAASP sampling has been implemented in BoTorch’s AF maximization and can optionally be enabled with the parameter sample_around_best. BoTorch augments a set of globally sampled candidates with the RAASP samples, resulting in twice as many initial candidates. It perturbs the best-performing points by replacing the value of each dimension with a probability of $\min\left(1,\frac{20}{d}\right)$ by a value drawn from a Gaussian distribution, truncated to stay within the bounds of the search space. BoTorch then chooses the points of maximum initial acquisition value to start the GD optimization of the AF.
With increasing dimensionality or descending length scale, the starting points for the multi-start GD routine chosen by the AF maximizer are increasingly more likely to originate from the RAASP samples. Fig. 4 (right panel) illustrates this. Here, we draw realizations from GPs, initialized with different dimensionalities ($x$-axis) and length scales ($y$-axis). For each realization, we maximize the AF with RAASP sampling and plot the percentage of candidates of maximum acquisition value originating from the RAASP samples across all candidates. A higher percentage indicates a more ‘local’ sampling. We further average across five random restarts. The percentage of RAASP candidates with maximum acquisition value increases with the input dimensionality and decreases with the length scale.
@@
In general, when the length scales of the GP are short and the dimensionality is high, BO shows a local-search-like behavior with RAASP sampling and a random search behavior without it. We demonstrate this using the observation traveling salesperson distance (OTSD(Papenmeier et al., 2025), which quantifies an algorithm’s exploration by finding the shortest path connecting all observations made up to a certain iteration. The OTSD curve of an algorithm $A$ consistently lying above the curve of algorithm $B$ indicates that algorithm $A$ is more explorative as its observations are spread more evenly across $\mathcal{X}$. The OTSD is always monotonic increasing. Fig. 5 shows the OTSDs for $100$-dimensional GPs with different length scales, each initialized with $10$ random samples in the design of experiments (DOE) phase and subsequently optimized with LogEI and RAASP sampling for $20$ iterations. Unless the model length scale is sufficiently long for the AF gradient not to vanish (as for $\ell=0.5$ in the right panel of Fig. 5), the AF maximizer picks one of the initial random candidates without further optimizing it. This is supported by the trajectories for the BO phase (iteration $\geq 9$) following the trajectory of the DOE phase (iteration $<9$) for $\ell=0.05$ and $\ell=0.28$ in the right panel of Fig. 5. We generally recommend the RAASP sampling method as it improves BO by automatically reverting to local search when encountering flat AFs.
@@ 4 Discussion
We propose a simple initialization for the gradient-based optimizer used for fitting the length scales of the GP surrogate via MLE and evaluate its performance for BO tasks. In what follows, we suppose a BO algorithm with a $\nicefrac{{5}}{{2}}$-ARD-Matérn kernel and LogEI (Ament et al., 2024). To address the issue of vanishing gradients at the start of the MLE optimization, we choose the initial length scale as 0.1 and scale with $\sqrt{d}$ to account for the increasing distances of the randomly sampled DOE points. Thus, the initial length scale used in the optimization is $\frac{\sqrt{d}}{10}$, and we refer to this new BO method as ‘MLE scaled’. The second method in the evaluation is the same BO method, but now using the default value $\ell=\ln 2$ of GPyTorch as initial length scale in MLE. This value is not scaled with $d$. Based on our analysis of the impact of RAASP sampling when optimizing the AF, we combine ‘MLE scaled’ with RAASP sampling and call the method ‘MLE Scaled with RAASP’ (MSR). We detail the RAASP sampling and the AF maximization in Appendix B. The next method, DSP, is the new default in BoTorch (Balandat et al., 2020), which uses a MAP estimate of length scales and initializes the optimization with the mode of the dimensionality-scaled length scale prior (DSP(Hvarfner et al., 2024). Table 1 summarizes the methods with basic BO setups we use for the empirical evaluation.
Table 1: Comparison of the simple BO methods used for the empirical evaluation.
Method length scale scaling RAASP sampling
MSR ✓(initial value)
MLE (scaled) ✓(initial value)
MLE ($\ell=\ln 2$)
DSP ✓(prior)
We also compare against SAASBO (Eriksson & Jankowiak, 2021), TuRBO (Eriksson et al., 2019), and Bounce (Papenmeier et al., 2023). SAASBO has a large computational runtime. Hence, we run it only for 500 iterations and terminate runs that exceed 72 hours. That is why SAASBO has fewer evaluations for Ant and Humanoid.
Our benchmarks are the $124$-dimensional soft-constrained version of the Mopta08 benchmark (Jones, 2008) introduced by Eriksson & Jankowiak (2021), the 180-dimensional Lasso-DNA (Šehić et al., 2022), the 388-dimensional SVM (Eriksson & Jankowiak, 2021), the 60-dimensional Rover (Eriksson et al., 2019), and two 888- and 6392-dimensional Mujoco benchmarks used by Hvarfner et al. (2024). The first four benchmarks are noise-free, while the others exhibit observational noise.
@@
Fig. 8 shows the performance of the BO methods on the four real-world applications. Each plot gives the average objective value of the best solution found so far in terms of the number of iterations. We show the ranking of the methods according to the final performance in Table 2 in Appendix C.1. The confidence bands indicate the standard error of the mean. MSR achieves competitive performance across all benchmarks, matching the SOTA DSP. Notably, MSR outperforms DSP on $124d$ Mopta08 and $888d$ Ant, and performs slightly worse than DSP other benchmarks. Bounce (Papenmeier et al., 2023) is consistently outperformed by MSR, MLE, and DSP but surpasses SAASBO (Eriksson & Jankowiak, 2021), especially on Ant. Although the constant length scale initialization ($\ell=\ln 2$) without RAASP achieves satisfactory results on lower-dimensional benchmarks such as $124d$ Mopta08 and $180d$ Lasso-DNA, it fails on higher-dimensional benchmarks like $888d$ Ant and $6392d$ Humanoid. We attribute this breakdown to vanishing gradients as shown in Fig. 13 in Appendix C.2.
@@
For the 124-dimensional Mopta08 and 180-dimensional Lasso-DNA applications, ‘MLE ($\ell=\ln 2$)’ learns length scales similar to the other methods. However, this constant initialization strategy fails to make progress from the initial value for the more high-dimensional problems in the bottom row of Fig. 9. We attribute this to vanishing gradients of the MLL as discussed in Sec. 3 and highlighted in Fig. 13 in Appendix C.2.
@@
Fig. 9 shows a surprising behavior for DSP. As one would anticipate, the estimated length scales are typically close to the mode of the hyperprior at the start of the optimization. However, they then converge to an even higher value on all benchmarks but the $6392$-dimensional Humanoid benchmark. Furthermore, the deviation from the prior mode is more pronounced for the lower-dimensional benchmarks, being in line with our analysis of the bias-variance trade-off in Sec. 3.3. At the beginning of its execution, the BO algorithm that uses MLE with scaled initial length scales (‘MLE (scaled)’) uses longer length scales than all other methods. The resulting estimates vary significantly for the high-dimensional Ant and Humanoid problems, supporting our analysis in Sec. 3.3 where we study the comparatively high variance of MLE compared to MAP.
@@
In Fig. 9, all methods converge to similarly long length scales on the Mopta08 and Lasso-DNA benchmarks. This is likely attributable to a specific property of these popular benchmarks, as we discuss in Appendix D. In short, these benchmarks have a simple structure that benefits models with long length scales, posing the risk of algorithms being ‘overfitted’ to these benchmarks.
@@ 5 Conclusions and Future Work
Our analysis reveals underlying challenges in HDBO while offering practical insights for improving HDBO methods. We demonstrate that common approaches for fitting GPs cause vanishing gradients in high dimensions. We propose an initialization for MLE that achieves state-of-the-art performance on established HDBO benchmarks without requiring the assumptions of maximum a-posteriori estimation. Finally, we provide empirical evidence that combining MLE with RAASP sampling reduces the variance of the length scale estimates and yields values closer to the ones learned by DSP, providing a fresh argument for the inclusion of RAASP sampling in HDBO.
In future work, we will continue to carefully vet popular benchmarks and propose novel, challenging benchmarks that preserve the traits of real-world applications. Furthermore, this work emphasizes the importance of numerical stability for the performance of BO in high dimensions. Thus, we propose approaching the development of models and AFs from this perspective.
Our work focuses on GP surrogate models but we will explore in how far our findings can be extended to other surrogate models, such as random forests or Bayesian neural networks.
Finally, we will explore how our findings help improve the performance of established techniques for HDBO by combining MSR with TRs (Eriksson et al., 2019), adaptive subspace embeddings (Papenmeier et al., 2022), or additive structures (Duvenaud et al., 2011).
Impact Statement